A gas mixture contained 42.18% by weight Xe and 57.82% Ne. What was the mole fraction of Xe?
Answer #1

Suppose, total weight of mixture = 100 gram

Then, mass of Xe = 42.18 gm

mass of Ne = 57.82 gm

We know, moles = mass/ molar mass

So, moles of Xe = 42.18/131.3 = 0.32 moles

moles of New = 57.82/20.2 = 2.86 moles

So, mole fraction of Xe

= moles of Xe/ total moles of Xe and Ne

= 0.32/0.32+2.86 = 0.32/3.18

= 0.10 (Answer)